OpenStudy (anonymous):

What makes the statement true? 8^-2=?

OpenStudy (anonymous):

.o156?

OpenStudy (anonymous):

Remember that \[x^{-n} = \frac{1}{x^n}\]

OpenStudy (anonymous):

.0156

OpenStudy (anonymous):

Recall: \[x^{-n} = \frac{ 1 }{x^{n} }\] Solution: \[=\frac{ 1 }{ 8^{2} } =\frac{ 1 }{ 64 } \approx 0.0156 \] So, yes.

OpenStudy (anonymous):

thank you!